wihtout solving, decide which meathod you would use to solve each system: graphing, substitution, or elimination. explain.
y=3x-1
y=4x
and
4s-3t=8
t=-2s-1
thanks for the help :)
appriciated

Answer by stanbon(75887)

wihtout solving, decide which meathod you would use to solve each system: graphing, substitution, or elimination. explain.
y=3x-1
y=4x
--------
Substitute for "y" and solve for "x":
4x = 3x-1
x = -1
---
Solve for "y":
y = 4x
y = 4*-1 = -4
----------------------------------
and
4s-3t=8
t=-2s-1
---
Substitute for "t" and solve for "s":
4s - 3(-2s-1) = 8
4s - 6s + 3 = 8
-2s = 5
s = -5/2
------------------------------------
Solve for "t":
t = -2s-1
t = -2(-5/2) - 1
--
t = 5-1 = 4
=================================
Cheers,
Stan H.